OpenStudy (anonymous):

From a rectangular sheet of sides 30cm and 40cm, a circular sheet as big as possible is cut off.Find the area of remaining sheet.

OpenStudy (shubhamsrg):

 owing to symmetry, radius of circle = 15 area = pi(15^2) so rem area = (30)(40) - pi(15^2)